# A Phase-field Model for Apoptotic Cell Death

### v2

@@ 3. Numerical Simulations
The model presented in Equations (12) is capable of simulating different types of topological transitions underlying the apoptotic phenomenon occurring in cells when induced by the presence of an apoptotic phase field. Here, we simulate specific characteristics of apoptotic processes, namely finger formation, fragmentation, and nucleation as seen in the work by You et al. [you2015kml001], with the aim of understanding parameter values triggering different transitions.
@@ 3.1. Physical and Numeric Parameter list
Table 1. Physical and chemical parameters required for simulating apoptosis. The parameters $\ell_{\sigma}^{2}$, $\ell_{\varphi}$, $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$ and $k_{2}$ are explored, and example values have been proposed in this table with the corresponding simulation for illustration.
@@
$k_{2}$ 8.0 (Figure 3), 1.0 (Figure 6) Parameter required for recreating reaction dynamics
@@
In this piece of work, we want to observe the effect on the morphological transitions of the cell during apoptosis. Thus, we vary and explore the reaction rate parameters $k_{1}$ and $k_{2}$, and thus vary the dimensionless parameters $\ell^{\varphi}_{r}$ and $\ell^{\sigma}_{r}$. Additionally, we will explore the parameter $\beta$, found in the dimensionless parameter $\ell_{2}$. However, as $\dot{\varphi}$ decreases, this will act as a sink term for $\sigma$, thus increasing the amount of consumption. The annihilation rate refers to the rate by which the cyto phase field is degraded by the apoptotic phase field. The rate of reaction is the rate at which the cell consumes the apoptotic phase field.
@@ 3.2. Initial conditions
While cellular shape can vary depending on cell type and environment [zeng2022cell], we restricted the cyto and apoptotic phase fields to circular shapes, with the addition of noise to mimic cellular inhomogeneities. In what follows, we define the initial conditions and noise. The noise standard deviation is $a=\sqrt{0.1}$ and normally distributed as $\eta(x,y)\sim\mathcal{N}(0,a^{2})$.
@@ 3.3. Finger formation and nucleation during apoptosis
In Equation (12), we describe the apoptotic degradation of the cyto phase field $\varphi$ while the apoptotic phase field $\sigma$ initiates the degradation. Figure 2 provides simulations illustrating the cyto phase field degradation when the interface is selected to be $\ell_{\varphi}=1.33\times 10^{-4}$ and the remaining parameters are $\ell_{\sigma}^{2}=1.5$, $\ell_{r}^{\varphi}=960$, $\ell_{r}^{\sigma}=0.28$ and $k_{2}=10.0$. This value of $\epsilon$ is chosen to produce a sharp interface based on the length scales for the size of the cyto phase field. In the simulation output, the cyto phase field $\sigma$ diminishes at a high rate, producing the formation of distinct fingers, at the expense of the apoptotic phase field $\varphi$. It can also be seen that nucleations can form in the cyto phase field. Based on the proposed model for the type of reaction included in Equations (2), degradation of the phase field $\varphi$ is promoted at the interface, that is, when $0<\varphi<1$. Therefore, inhomogeneities in the initial conditions allow for reactions to occur in the interior of the cyto phase field, hence enabling nucleations to form. However, these nucleations happen at a much lower rate. For this reason, it can be seen that the apoptotic phase field is consumed at the cyto phase field interface (membrane), starting at $\bar{t}$=3.98 and continuing in $\bar{t}$=7.99 and $\bar{t}$=11.76.
@@
Figure 2. Progression of the $\varphi$ and $\sigma$ phase fields at the following nondimensionalised time points at $\bar{t}=0.00\ \text{(A)},\ 3.98\ \text{(B)},\ 7.99\ \text{(C)},\ 11.76\ \text{(D)}$. Using parameter values of: $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$ and $k_{2}=10.0$. The remaining parameters are listed in Table 1.
@@ 3.4. Interface width influence
Figure 3. Evolution at nondimensionalised time $\bar{t}=0.00$, (A) 4.14 (B), 8.35 (C), 11.98 (D) of the phase field $\varphi$ with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. Distinct finger formation can be observed.
@@
Figure 4. Evolution at $\bar{t}=0$ (A), 4.08 (B), 7.51 (C), 11.95 (D) of $\varphi$ with interface width increased to $\ell_{\varphi}=5.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$ and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. Finger formation in $\varphi(\boldsymbol{x},t)$ can be observed as forming during the degradation of the phase field. A fragment can be seen to form from one finger.
@@
Figure 5. Evolution at nondimensionalised time $\bar{t}=0.00$ (A), 2.57 (B), 4.42 (C), 11.42 (D) of $\varphi$ with an interface width $\ell_{\varphi}=2.133\times 10^{-3}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. No finger formation observed around the interface and fragmentation of the phase field $\varphi(\boldsymbol{x},t)$ during apoptosis.
It is known in phase field models that changes to the interface width can have a range of effects on the resulting dynamics, as a broader interface leads to a loss of detail in the simulation, as stated by Qin and Bhadeshia [qin2010phase]. As can be seen in Figures 3, 4, and 5, a lower $\tilde{\epsilon}$ value sharpens the interface, while at larger values, the phase field has a diffuse interface. Figure 5 has a ratio of interface width to dynamic modulus of $\ell_{\varphi}=2.133\times 10^{-3}$, and here, there is no finger formation during the degradation and degradation of the phase field $\varphi$. However, by reducing $\ell_{\varphi}=5.333\times 10^{-4}$ (Figure 4), fingers begin to form during the degradation of the phase field $\varphi$. At $\ell_{\varphi}=1.333\times 10^{-4}$, finger formation in $\varphi$ is more distinct compared to simulations obtained with higher values for $\ell_{\varphi}$, as can be seen in Figures 4 and 5.
As seen in the bottom right-hand corner of Figure 4 (C) and in Figure 5 (B), (C), and (D), there is a pinch-off (fragmentation) from the finger during phase field $\varphi$ degradation. This effect occurs at a specific interface width and cannot be found when $\ell_{\varphi}=1.333\times 10^{-4}$. The pinch-off results from the larger surface area present due to a larger interface width, hence promoting a higher number of reactions and faster topological transition of the phase field $\varphi$. Here, we analyse two sets of reaction rate parameters, given by $(k_{1},k_{2})\in\{(0.9,10.0),\,(0.9,8.0),\,(4.0,1.0)\}$. The choice of these parameters influences the behaviour of $r(\varphi,\sigma)$ in Equation (13). When $k_{1}=4.0$ and $k_{2}=1.0$, the reaction term $r$ approximates $r\approx g(\varphi)\sigma$. In contrast, the parameter choices $(k_{1},k_{2})=(0.9,10.0)$ or $(0.9,8.0)$ yield an approximation of $r\approx g(\varphi)\sigma^{1/3}$. However, implementing $r$ directly in this latter form leads to numerical instabilities; therefore, the expression for $r$ as presented in Equation (13) is preferred.
@@ 3.5. Varying the rate of reaction and annihilation
3.5.1. Rate of apoptosis when varying the rate of reaction parameters
@@
Figure 6. Evolution at $\bar{t}=0.00$ (A), 1.03 (B), 3.87 (C), 11.97 (D) of the phase field $\varphi$ with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters have the values listed in Table 1. A higher reaction rate leads to the formation of a greater number of fragments when compared with Figure 3.
@@ 3.5.2. Varying the rate of annihilation for the cell
Figure 7. Evolution at $\bar{t}=0.00$ (A), 2.88 (D), 4.32 (C), 5.28 (D) of the phase field $\varphi$ with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=2.0$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters are listed in Table 1. Larger fragments form when $\beta=1.5$ and no nucleation.
@@
Figure 8. Evolution at $\bar{t}=0.00$ (A), 2.26 (B), 4.37 (C), 5.98 (D) of the phase field $\varphi$ with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=5.0$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters are listed in Table 1. Nucleation is increased, and fragmentation is decreased compared to $\beta=1.5$ in Figure 6.
@@ 3.6. Internal configurational forces
Here, we compute the internal configurational force, given in Equation (22) in Appendix A, to analyse the driving forces triggering the different topological transitions described in previous sections. We restrict attention to regions surrounding the cyto phase field and nucleations, as depicted in Figure 2.
@@
Figure 9. Evolution of a section of the interface for the phase fields $\varphi$ and $\sigma$, and the internal configurational forces $\boldsymbol{f}$, and $|\boldsymbol{f}|$ at $\bar{t}$=0.19 (A, E(i)), 1.20 (B, E(ii)), 2.64 (C, E(iii)), and 3.84 (D, E(iv)). The parameters selected for this simulation as listed in Table 1, $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The arrows in images (A), (B), (C), and (D) indicate the direction of $\boldsymbol{f}$; their size reflects the relative magnitude, while the colour represents the magnitude throughout the simulation—red denotes higher values than blue. In E(i-iv), the $|\boldsymbol{f}|$ is overlayed with the phase-field $\varphi(\boldsymbol{x},t)$ with red regions indicating a greater magnitude and blue regions indicating a lower magnitude.
Figure 9 shows details of the interface encompassing the cyto phase field $\varphi$ and the apoptotic phase field $\sigma$ with arrows representing the internal configurational forces. At $\bar{t}=0.19$, in Figure 9 (A), the internal configurational force $\boldsymbol{f}$ is nearly uniform in space along the interface. Therefore, the interfaces for the phase fields $\varphi$ and $\sigma$ degrade uniformly. As time progresses to $\bar{t}=1.20$, Figure 9 (B), the internal configurational oscillates along the interface. Thus, it produces the early stages of finger formation. At $\bar{t}=2.64\ \text{and}\ 3.84$, Figures 9 (C) and (D), distinct fingers form in the phase field $\varphi$. Furthermore, the magnitude of the internal configurational forces increases at the basin of the opening that forms between two fingers.
@@
Figure 10. Evolution of a nucleation in $\varphi$ and $\sigma$. The internal configurational force $\boldsymbol{f}$ is present at the nucleation interface. The nucleation is at $\bar{t}$=0.41 (A, E(i)), 1.44 (B, E(ii)), 2.64 (C, E(iii)), and 3.84 (D, E(iv)).The parameters selected for this simulation as listed in Table 1, $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The arrows in images (A), (B), (C), and (D) indicate the direction of $\boldsymbol{f}$; their size reflects the relative magnitude, while the colour represents the magnitude throughout the simulation—red denotes higher values than blue. In E(i-iv), the $|\boldsymbol{f}|$ is overlayed with the phase-field $\varphi(\boldsymbol{x},t)$ with red regions indicating a greater magnitude and blue regions indicating a lower magnitude.
[…]
4. Qualitative comparison of electron microscopy images and simulations

### v3

@@ 3. Numerical Simulations
The model presented in Equations (15) is capable of simulating different types of topological transitions underlying the apoptotic phenomenon occurring in cells when induced by the presence of an cytotoxic phase field. Here, we simulate specific characteristics of apoptotic processes, namely finger formation, fragmentation, and nucleation as seen in the work by You et al. [25], with the aim of understanding parameter values triggering different transitions.
@@ 3.1. Physical and Numeric Parameter list
Table 1. Physical and chemical parameters required for simulating apoptosis. The parameters $\ell_{\sigma}^{2}$, $\ell_{\varphi}$, $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$ and $k_{2}$ are explored, and example values have been proposed in this table with the corresponding simulation for illustration. Remaining parameters are fixed at their values presented in the table in all simulations.
@@
$k_{2}$ 8.0 (Figure 3), 1.0 (Figure 6) Sensitivity to anisotropy
@@
In this piece of work, we want to observe the effect on the morphological transitions of the cell during apoptosis. Thus, we vary and explore the reaction rate parameters $k_{1}$ ($\ell^{\sigma}_{r}$) and $k_{2}$, and thus vary the dimensionless parameters $\ell^{\varphi}_{r}$ and $\ell^{\sigma}_{r}$. Additionally, we will explore the parameter $\beta$, found in the dimensionless parameter $\ell_{2}$. As the parameter $\beta$ increases, this will act as a greater sink term for $\sigma$, thus increasing the amount of consumption. The annihilation rate refers to the rate by which the cyto phase field is degraded by the cytotoxic phase field. The rate of reaction is the rate at which the cell consumes the cytotoxic phase field.
@@ 3.2. Initial conditions
While cell shape can vary depending on cell type and environment [26], we restricted the cyto and cytotoxic phase fields to circular shapes, with the addition of noise to mimic cellular inhomogeneities. In what follows, we define the initial conditions and noise. The noise standard deviation is $a=\sqrt{0.1}$ and normally distributed as $\eta(x,y)\sim\mathcal{N}(0,a^{2})$.
@@ 3.3. Finger formation and nucleation during apoptosis
In Equation (15), we describe the apoptotic degradation of the cyto phase field $\varphi$ while the cytotoxic phase field $\sigma$ initiates the degradation. Figure 2 provides simulations illustrating the cyto phase field degradation when the interface is selected to be $\ell_{\varphi}=1.33\times 10^{-4}$ and the remaining parameters are $\ell_{\sigma}^{2}=1.5$, $\ell_{r}^{\varphi}=960$, $\ell_{r}^{\sigma}=0.28$ and $k_{2}=10.0$. This value of $\epsilon$ is chosen to produce a sharp interface based on the length scales for the size of the cyto phase field. In the simulation output, the cyto phase field $\varphi$ diminishes at a high rate, producing the formation of distinct fingers, at the expense of the cytotoxic phase field $\sigma$. It can also be seen that nucleations can form in the cyto phase field. Based on the proposed model for the type of reaction included in Equations (2), degradation of the phase field $\varphi$ is promoted at the interface, that is, when $0<\varphi<1$. Therefore, inhomogeneities in the initial conditions allow for reactions to occur in the interior of the cyto phase field, hence enabling nucleations to form. However, these nucleations happen at a much lower rate. For this reason, it can be seen that the cytotoxic phase field is consumed at the cyto phase field interface (membrane), starting at $\bar{t}$=3.98 and continuing in $\bar{t}$=7.99 and $\bar{t}$=11.76.
@@
Figure 2. Progression of the $\varphi$ and $\sigma$ phase fields at the following nondimensionalised time points at $\bar{t}=0.00\ \text{(A)},\ 3.98\ \text{(B)},\ 7.99\ \text{(C)},\ 11.76\ \text{(D)}$. Using parameter values of: $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$ and $k_{2}=10.0$. The remaining parameters in the simulation are listed in Table 1.
@@ 3.4. Interface width influence
Figure 3. Evolution of the phase field $\varphi$ at the nondimensionalised time points $\bar{t}=0.00$, (A) 4.14 (B), 8.35 (C), 11.98 (D) with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. Distinct finger formation can be observed.
@@
Figure 4. Evolution of the phase field $\varphi$ at the time points at $\bar{t}=0$ (A), 4.08 (B), 7.51 (C), 11.95 (D) of $\varphi$ with interface width increased to $\ell_{\varphi}=5.333\times 10^{-4}$ in comparison to the simulation presented in Figure 3. The following parameters $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$ and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. Finger formation in $\varphi(\boldsymbol{x},t)$ can be observed as forming during the degradation of the phase field. A fragment can be seen to form from one finger.
@@
Figure 5. Evolution of the phase field $\varphi$ at nondimensionalised time $\bar{t}=0.00$ (A), 2.57 (B), 4.42 (C), 11.42 (D) of $\varphi$ with an interface width $\ell_{\varphi}=2.133\times 10^{-3}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The remaining parameters required to recreate the simulation are as listed in Table 1. No finger formation observed around the interface and fragmentation of the phase field $\varphi(\boldsymbol{x},t)$ is present during this simulation of apoptosis.
It is known in phase-field equations that changes to the interface width can have a range of effects on the resulting dynamics, as a broader interface leads to a loss of detail in the simulation, as stated by Qin and Bhadeshia [22]. As can be seen in Figures 3, 4, and 5, a lower $\tilde{\epsilon}$ value sharpens the interface, while at larger values, the phase field has a diffuse interface. Figure 5 has a ratio of interface width to dynamic modulus of $\ell_{\varphi}=2.133\times 10^{-3}$, and here, there is no finger formation during the degradation of the phase field $\varphi$. However, by reducing $\ell_{\varphi}=5.333\times 10^{-4}$ (Figure 4), fingers begin to form during the degradation of the phase field $\varphi$. At $\ell_{\varphi}=1.333\times 10^{-4}$, finger formation in $\varphi$ is more distinct compared to simulations obtained with higher values for $\ell_{\varphi}$, as can be seen in Figures 4 and 5.
As seen in the bottom right-hand corner of Figure 4 (C) and in Figure 5 (B), (C), and (D), there is a pinch-off (fragmentation) from the finger during phase field $\varphi$ degradation. This event occurs at a specific interface width and cannot be found when $\ell_{\varphi}=1.333\times 10^{-4}$. The pinch-off results from the larger surface area present due to a larger interface width, hence promoting a higher number of reactions and faster topological transition of the phase field $\varphi$. Here, we analyse two sets of reaction rate parameters, given by $(k_{1},k_{2})\in\{(0.9,10.0),\,(0.9,8.0),\,(4.0,1.0)\}$. The choice of these parameters influences the behaviour of $r(\varphi,\sigma)$. When $k_{1}=4.0$ and $k_{2}=1.0$, the reaction term $r$ approximates $r\approx g(\varphi)\sigma$. In contrast, the parameter choices $(k_{1},k_{2})=(0.9,10.0)$ or $(0.9,8.0)$ yield an approximation of $r\approx g(\varphi)\sigma^{1/3}$. However, implementing $r$ directly in this latter form leads to numerical instabilities; therefore, the expression for $r$ as presented in the system of Equations (12) is preferred.
@@ 3.5. Varying the rate of reaction and annihilation
3.5.1. Rate of apoptosis when varying of reaction parameters
@@
Figure 6. Evolution of the phase field $\varphi$ at the time points $\bar{t}=0.00$ (A), 1.03 (B), 3.87 (C), 11.97 (D) with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters have the values listed in Table 1. A higher reaction rate leads to the formation of a greater number of fragments when compared with Figure 3.
@@ 3.5.2. Varying the rate of annihilation for the cell
Figure 7. Evolution of the phase field $\varphi$ at the time points $\bar{t}=0.00$ (A), 2.88 (D), 4.32 (C), 5.28 (D) with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=2.0$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters are listed in Table 1. Larger fragments form when $\beta=1.5$ and no nucleation.
@@
Figure 8. Evolution of the phase field $\varphi$ at the time points $\bar{t}=0.00$ (A), 2.26 (B), 4.37 (C), 5.98 (D) of the phase field $\varphi$ with $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=5.0$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$. The remaining parameters are listed in the Table 1. Nucleation is increased, and fragmentation is decreased compared to $\beta=1.5$ in Figure 6.
@@ 3.6. Internal configurational forces
In the context of apoptosis, the loss of the actin cytoskeleton and other structural components removes the support that maintains cellular shape, allowing the cell to explore a new configurational space. As a result, the dominant driving mechanisms become interfacial, with configurational forces acting to reorganise the cell boundary and internal interfaces. Here, we compute the internal configurational force, given in Equation (25) in the Appendix S A, to analyse the driving forces triggering the different topological transitions described in the previous section. We restrict attention to regions surrounding the cyto phase field and nucleations, as depicted in Figure 2.
@@
Figure 9. Evolution of a section of the interface for the phase fields $\varphi$ and $\sigma$, and the internal configurational forces $\boldsymbol{f}$, and $|\boldsymbol{f}|$ at the time points $\bar{t}$=0.19 (A, E(i)), 1.20 (B, E(ii)), 2.64 (C, E(iii)), and 3.84 (D, E(iv)). The parameters selected for this simulation as listed in Table 1, $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The arrows in images (A), (B), (C), and (D) indicate the direction of $\boldsymbol{f}$; their size reflects the relative magnitude, while the colour represents the magnitude throughout the simulation—red denotes higher values than blue. In E(i-iv), the evolution of the phase fields $\varphi$ and $\sigma$ have been plotted across the domain, illustrating the global degradation of these phases.
Figure 9 shows details of the interface encompassing the cyto phase field $\varphi$ and the cytotoxic phase field $\sigma$ with arrows representing the internal configurational forces. At $\bar{t}=0.19$, in Figure 9 (A), the internal configurational force $\boldsymbol{f}$ is nearly uniform in space along the interface. Therefore, the interfaces for the phase fields $\varphi$ and $\sigma$ degrade uniformly. As time progresses to $\bar{t}=1.20$, Figure 9 (B), the internal configurational oscillates along the interface. Thus, it produces the early stages of finger formation. At $\bar{t}=2.64\ \text{and}\ 3.84$, Figures 9 (C) and (D), distinct fingers form in the phase field $\varphi$. Furthermore, the magnitude of the internal configurational forces increases at the basin of the opening that forms between two fingers.
@@
Figure 10. Evolution of a nucleation in $\varphi$ and $\sigma$. The internal configurational force $\boldsymbol{f}$ is present at the nucleation interface. The nucleation is at $\bar{t}$=0.41 (A, E(i)), 1.44 (B, E(ii)), 2.64 (C, E(iii)), and 3.84 (D, E(iv)).The parameters selected for this simulation as listed in Table 1, $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$. The arrows in images (A), (B), (C), and (D) indicate the direction of $\boldsymbol{f}$; their size reflects the relative magnitude, while the colour represents the magnitude throughout the simulation (red denotes higher values than blue). In E(i-iv), the evolution of the phase fields $\varphi$ and $\sigma$ have been plotted across the domain, illustrating the global degradation of these phases.
Figure 10 depicts the development of internal configurational forces during nucleation as the phase fields $\varphi$ and $\sigma$ degrade. At $\bar{t}=0.41$, Figure 10 (A), the internal configurational force has a greater magnitude at the nucleation interface than at the main interface surrounding the phase field $\varphi$. The same behaviour can be observed in the phase field $\sigma$. Even though the nucleation appears to be circular, in Figure 10 (A), the magnitude of the internal configurational force along the interface surrounding the nucleation is not uniform. Despite the disuniformity in the internal configurational force magnitude, the cavity expands rapidly in all directions. As the simulation progresses to $\bar{t}$=1.44 (Figure 10 (B)), 2.64 (Figure 10 (C)), and 3.84 (Figure 10 (D)), the internal configuration force disappears at the nucleation interface closest to the interface encompassing the bulk of phase fields $\varphi$ and $\sigma$. Therefore, the reduction in the degradation of the phase field $\varphi$ results from the phase field $\sigma$ no longer being present in the space between the nucleation and the main interface. At the later time points $\bar{t}$=1.44 (Figure 10 (B)), 2.64 (Figure 10 (C)), and 3.84 (Figure 10 (D)), the cytotoxic phase field $\sigma$ has reacted with the cyto phase field $\varphi$ and been completely depleted. After the cytotoxic phase field is consumed, the internal configurational force vanishes, and the degradation in the cyto phase field $\varphi$ ceases. In the lower area of the nucleation at $\bar{t}=3.84$ (Figure 10 (D)), it is observed that the magnitude of the internal configurational forces develops in two distinct regions on either side of a finger forming.
4. Comparison of electron microscopy images and simulations